# GIVE: Structured Reasoning of Large Language Models with Knowledge Graph Inspired Veracity Extrapolation

### v2

@@ Abstract
Existing approaches based on context prompting or reinforcement learning (RL) to improve the reasoning capacities of large language models (LLMs) depend on the LLMs’ internal knowledge to produce reliable Chain-Of-Thought (CoT). However, no matter the size of LLMs, certain problems cannot be resolved in a single forward pass. Meanwhile, agent-based reasoning systems require access to a comprehensive nonparametric knowledge base, which is often costly or not feasible for use in scientific and niche domains. We present Graph Inspired Veracity Extrapolation (GIVE), a novel reasoning method that merges parametric and non-parametric memories to improve accurate reasoning with minimal external input. GIVE guides the LLM agent to select the most pertinent expert data (observe), engage in query-specific divergent thinking (reflect), and then synthesize this information to produce the final output (speak). Extensive experiments demonstrated the following benefits of our framework: (1) GIVE boosts the performance of LLMs across various sizes. (2) In some scenarios, GIVE allows smaller LLMs to surpass larger, more sophisticated ones in scientific tasks (GPT3.5T + GIVE GPT4). (3) GIVE is effective on scientific and open-domain assessments. (4) GIVE is a training-free method that enables LLMs to tackle new problems that extend beyond their training data (up to 43.5% 88.2% accuracy improvement). (5) GIVE allows LLM agents to reason using both restricted (very small) and noisy (very large) knowledge sources, accommodating knowledge graphs (KG) ranging from 135 to more than 840k nodes. (6) The reasoning process involved in GIVE is fully interpretable.
@@ 1 Introduction
Context-based methods (Wei et al., 2023; Brown et al., 2020) to enhance the reasoning ability of large language models (LLMs) (Ouyang et al., 2022; Chowdhery et al., 2022; OpenAI et al., 2024b; Grattafiori et al., 2024; OpenAI et al., 2024a) incorporate examples with logic chains in the prompt, allowing the model to generate analogous logical sequences for the given query. Recent investigations (Zelikman et al., 2024; DeepSeek-AI et al., 2025; Shao et al., 2024) highlighted the substantial promise of reinforcement learning (RL) to enhance the generation of high-quality logic. These methods operate on the premise that parametric memory of LLM is sufficient to execute accurate reasoning. They demonstrate satisfactory performance in reasoning-complex tasks that do not require novel knowledge, such as mathematical and commonsense. However, they fail to achieve comparable enhancements in scientific tasks due to the insufficiency of pre-trained knowledge (Cai et al., 2024; Zhang et al., 2024; Dorfner et al., 2024; Dong et al., 2023; Deng et al., 2024). Such information is rare in the web data, in fact, irrespective of the number of parameters, LLMs face unsolvable challenges using a single forward pass (Dziri et al., 2024). Incorporating external knowledge becomes necessary to enable the model to adapt to tasks beyond the training set. A line of research (Sun et al., 2024; Luo et al., 2024; Xu et al., 2024) advocates the iterative exploration of a knowledge base, and advanced retrieval-augmented generation (RAG) frameworks (Mavromatis & Karypis, 2024; Edge et al., 2024) model the interconnections between information units to improve knowledge coherence. These strategies have proven effective in specific question answering scenarios, assuming the existence of an accessible and retrievable database containing the correct reasoning logic. However, in scientific realms, the construction of comprehensive knowledge bases is daunting, which requires progress in both domain-specific natural language processing (NLP) and field-wide vocabulary standardization (Badal et al., 2019; Verhagen et al., 2012). Enhancing LLM reasoning with limited external information is a more realistic approach. Although the retrieved knowledge lacks direct evidence for problem solving, it embodies the intuition and expertise of curation specialists, such as feasible relation collection and potential links among similar entities.
@@
In this paper, we seek to overcome the constraints of relying solely on either internal or external knowledge by introducing GIVE, a graph-inspired veracity extrapolation framework. GIVE emulates the cognitive processes employed by human experts, draws inspiration from related knowledge, and conducts divergent reasoning. It populates the limited information with provisional connections between query concepts, informed by factual linkages, and grounds these links using LLMs’ parametric knowledge. Our method also constructs counterfactual reasoning to mitigate hallucinations and incorporates intermediate entities for multi-step reasoning. Specifically, GIVE identifies a concentrated group of entities closely associated with the query. By exploring potential relations within the pertinent knowledge graph (KG) concepts, we develop a reasoning framework that encompasses all conceivable concepts and connections that could enhance query resolution. We incorporate additional intermediate concepts by selecting multi-step reasoning strategies that most effectively address the questions. GIVE comprises retrieved expert information, internal knowledge acquired through pre-training, and innovative relations that unite analogous concepts via veracity extrapolation. To complete the reasoning framework, counterfactual linkages among nodes are integrated to avert hallucinations. Ultimately, GIVE is a structured reasoning framework of LLMs that (1) retrieves external knowledge to increase the informativeness of responses; (2) conduct divergent reasoning on the limited expert information that does not directly solve the query, by extrapolating expert triplets to correlated queried concepts, termed "veracity extrapolation." We test our proposed approach on both scientific and open-domain tasks, including biomedical, commonsense, and open-domain question answering. GIVE consistently attains superior performance across all datasets, utilizing KGs of differing sizes and densities, surpassing all internal/external knowledge reasoning benchmarks, which demonstrates the effectiveness and reliability of the proposed framework. Encouragingly, GIVE enables smaller LLMs such as GPT3.5T to achieve better performance than advanced models like GPT4 in scientific tasks. GIVE pioneers in elevating LLM’s reasoning capabilities with minimal external cues to activate its intrinsic problem-solving capacity.
@@ 2 Problem Statement
Assume that represents the accurate logic chain for solving query . Reasoning on internal knowledge techniques (Wei et al., 2023; Brown et al., 2020; DeepSeek-AI et al., 2025) make use of reinforcement learning or context-based prompts to direct the LLM to produce an accurate logic chain with high probability, relying solely on the model’s intrinsic knowledge. This signifies dependence on parameter , such that . In contrast, reasoning with retrieval frameworks (Sun et al., 2024), posits the availability of a comprehensive knowledge base , and employs a retrieval model such that .
@@
While theoretically, this offers no benefit over RAG (Lewis et al., 2021) for generating answers with retrieved context, numerous studies (Ge et al., 2023; Bang et al., 2023) have shown that LLMs lack expertise in scientific domains and cannot construct a multi-step logical chain (Wei et al., 2023; Wang et al., 2023; Jiang et al., 2024) to connect a complex query with the incomplete information that cannot directly solve it. GIVE, in this context, provides additional guidance to LLMs, helping them to populate the expert knowledge towards the key components of the query by integrating both parametric and non-parametric data.
@@ 3.2 Entity Group Construction
,The set includes entities semantically akin to , and is added to to create the entity group :
@@ 3.3 Inner-group connections
Initially, focusing on each in Section 3.2, we establish links between the queried entity and its semantically-related concepts within its entity group. This aims to prompt the LLM to explore related concepts broadly, rather than focusing solely on the queried entity. The challenge of directly inducing relationships between two queried entities is addressed by identifying possible links between two sets of similar concepts. We employ the LLM to suggest relationships between the queried entity and each in-group concept. All such knowledge are appended to an affirmative knowledge set. Consider an entity group with 1 queried entity and additional KG concepts , for :
@@ 3.4.1 Potential Relations Induction
We first determine all possible relationships that could link any node pairs within the two groups. We evaluate two categories of potential relations: (1) Relations specified in the question. These are the vital connections required for the ultimate QA task. (2) KG relations present between these groups. Due to the semantic similarity of concepts within each group, the existing cross-group KG connections may link other entity pairs across groups. Formally, for KG , entity groups and , we define the set of potential relations as:
@@ 3.4.2 Intermediate Node Group Discovery for Multi-step Reasoning
In scientific fields, directly establishing links between specific concepts is sometimes not feasible. For instance, to address a query regarding a drug’s effect on a disease, a logical approach is to associate these through (drug, compound, disease) links, forming the statement: "since a certain compound exists within the drug, and this compound alleviates particular diseases, it can be inferred that the drug may treat the disease." To support such multi-step reasoning process, GIVE investigates novel node groupings used as intermediate steps in reasoning. GIVE directs LLM to identify the most beneficial 2-hop paths between groups, aiding the question-answering task. Through the intermediate node of an optimal 2-hop path, GIVE forms an intermediate entity group using the process outlined in Section 3.2.
@@ 3.4.3 Structured expert knowledge guided reasoning
Finally, GIVE introduces divergent thinking on these non-parametric evidence, including concept groups and all potential relations between them.
Relation Assignment Using External Evidence. If an edge exists on the external knowledge graph between two entities, we inherit the ground-truth relation from the original KG. When verbalizing such edges, we use a definite tense to convey the high certainty of this fact.
@@
Identifying Open Relations for novel Connections. To avoid the limitations of the knowledge base’s scope, GIVE also offers LLM the freedom to independently generate a short phrase describing the relation of each entity pair.
@@ 3.5 Progressive answer generation
From the reasoning process presented in Section 3.4.3, GIVE obtained three kinds of knowledge: (1) Affirmative knowledge set that contains all knowledge affirmed by LLMs’ parametric knowledge, including inner-group connections; the cross-group open connections, and the extrapolated connections. We refer to this knowledge set as . (2) Counter-factual knowledge set that contains all rejected potential cross-group connections, which is referred to as , (3) The ground-truth KG knowledge set . To prevent hallucination, we adopt a progressive manner to generate the final answer by first giving only the affirmative knowledge set. GIVE further directs LLM to refine this answer with the full context and the counter-factual knowledge set. The final answer is generated by providing all previous context and the ground-truth knowledge set. Given generator and the retrieved knowledge sets:
@@ 5 Conclusion
Acknowledgement
This effort was partially sponsored by NSF grants 2200274, 2106859 and 2312501, as well as NIH grants U54HG012517 and U24DK097771.
@@ B.1 Different ways of prompting
B.2 Number of seeded examples
B.3 Encoding model size
@@ C.1 Efficiency of GIVE
We further conduct experiments to compare the efficiency of GIVE with RAG (Lewis et al., 2021) and ToG (Sun et al., 2024), in terms of running time and context length, for every experiment setting we include in Section 4, the results are presented in Table 8.

### v3

# GIVE: Structured Reasoning of Large Language Models with Knowledge-Graph-Inspired Veracity Extrapolation

@@ Abstract
Existing approaches based on context prompting or reinforcement learning (RL) to improve the reasoning capacities of large language models (LLMs) depend on the LLMs’ internal knowledge to produce reliable Chain-Of-Thought (CoT). However, no matter the size of LLMs, certain problems cannot be resolved in a single forward pass. Meanwhile, agent-based reasoning systems require access to a comprehensive nonparametric knowledge base, which is often costly or not feasible for use in scientific and niche domains. We present Graph Inspired Veracity Extrapolation (GIVE), a novel reasoning method that merges parametric and non-parametric memories to improve accurate reasoning with minimal external input. GIVE guides the LLM agent to select the most pertinent expert data (observe), engage in query-specific associative thinking (reflect), and then synthesize this information to produce the final output (speak). Extensive experiments demonstrated the following benefits of our framework: (1) GIVE increases the performance of LLMs across various sizes. (2) In some scenarios, GIVE allows smaller LLMs to surpass larger, more sophisticated ones in scientific tasks (GPT3.5T + GIVE GPT4). (3) GIVE is effective on scientific and open-domain assessments. (4) GIVE is a training-free method that enables LLMs to tackle new problems that extend beyond their training data (up to 43.5% 88.2% accuracy improvement). (5) GIVE allows LLM agents to reason using both restricted (very small) and noisy (very large) knowledge sources, accommodating knowledge graphs (KG) ranging from 135 to more than 840k nodes. (6) The reasoning process involved in GIVE is fully interpretable. Our code is available at https://github.com/Jason-Tree/GIVE
@@ 1 Introduction
Context-based methods (Wei et al., 2023; Brown et al., 2020) to enhance the reasoning ability of large language models (LLMs) (Ouyang et al., 2022; Chowdhery et al., 2022; OpenAI et al., 2024b; Grattafiori et al., 2024; OpenAI et al., 2024a) incorporate examples with logic chains in the prompt, allowing the model to generate analogous logical sequences for the given query. Recent investigations (Zelikman et al., 2024; DeepSeek-AI et al., 2025; Shao et al., 2024) highlighted the substantial promise of reinforcement learning (RL) in improving the generation of high-quality logic. These methods operate on the premise that the parametric memory of LLM is sufficient to execute accurate reasoning. They demonstrate satisfactory performance in reasoning-complex tasks that do not require novel knowledge, such as mathematical and common sense. However, they fail to achieve comparable enhancements in scientific tasks due to the insufficiency of pre-trained knowledge (Cai et al., 2024; Zhang et al., 2024; Dorfner et al., 2024; Dong et al., 2023; Zhong et al., 2023). This type of information is rare in web data; in fact, regardless of the number of parameters, LLMs face unsolvable challenges using a single forward pass (Dziri et al., 2024). Incorporating external knowledge becomes necessary to enable the model to adapt to tasks beyond the training set.
Recent studies have highlighted the significant potential of incorporating structured knowledge bases (KGs) into LLM inference processes to enhance knowledge provision. Advanced retrieval-augmented generation (RAG) frameworks (Liang et al., 2024; Panda et al., 2024) identify precise KG from documents to aid accurate information retrieval for specific queries; research efforts are also proposed to promote iterative exploration (Sun et al., 2024; Luo et al., 2024; Xu et al., 2024) and improved modeling of interconnections within knowledge bases to improve coherence (Mavromatis & Karypis, 2024; Edge et al., 2024).
These strategies have proven effective in specific question answering scenarios, assuming the existence of an accessible and retrievable database containing the correct reasoning logic. However, in scientific realms, the construction of comprehensive knowledge bases is daunting, which requires progress in both domain-specific natural language processing (NLP) and field-wide vocabulary standardization (Badal et al., 2019; Verhagen et al., 2012). Enhancing LLM reasoning with limited external information is a more realistic approach. Although the retrieved knowledge lacks direct evidence for problem solving, it embodies the intuition and expertise of curation specialists, such as feasible relation collection and potential links among similar entities.
@@
In this paper, we seek to overcome the constraints of relying solely on either internal or external knowledge by introducing GIVE, a graph-inspired veracity extrapolation framework. GIVE emulates the cognitive processes employed by human experts, draws inspiration from related knowledge, and conducts associative thinking. It populates the limited information with provisional connections between query concepts, informed by factual linkages, and grounds these links using LLMs’ parametric knowledge. Our method also constructs counterfactual reasoning to mitigate hallucinations and incorporates intermediate entities for multi-step reasoning. Specifically, GIVE identifies a concentrated group of entities closely associated with the query. By exploring potential relations within pertinent knowledge graph (KG) concepts, we develop a reasoning framework that encompasses all conceivable concepts and connections that could enhance query resolution. We incorporate additional intermediate concepts by selecting multistep reasoning strategies that most effectively address the questions. GIVE comprises retrieved expert information, internal knowledge acquired through pre-training, and innovative relations that unite analogous concepts via veracity extrapolation. To complete the reasoning framework, counterfactual links among nodes are integrated to avert hallucinations. Ultimately, GIVE is a structured reasoning framework of LLMs that (1) retrieves external knowledge to increase the informativeness of responses; (2) conduct divergent reasoning on the limited expert information that does not directly solve the query, by extrapolating expert triplets to correlated queried concepts, termed "veracity extrapolation." We test our proposed approach on both scientific and open-domain tasks, including biomedical, commonsense, and open-domain question answering. GIVE consistently achieves superior performance across all datasets, utilizing KGs of differing sizes and densities, surpassing all internal/external knowledge reasoning benchmarks, which demonstrates the effectiveness and reliability of the proposed framework. Encouragingly, GIVE enables smaller LLMs such as GPT3.5T to achieve better performance than advanced models like GPT4 in scientific tasks. GIVE pioneers in elevating LLM’s reasoning capabilities with minimal external cues to activate its intrinsic problem-solving capacity.
@@ 2 Problem Statement
Assume that represents the accurate logic chain to solve the query . Reasoning on internal knowledge techniques (Wei et al., 2023; Brown et al., 2020; DeepSeek-AI et al., 2025) make use of reinforcement learning or context-based prompts to direct the LLM to produce an accurate logic chain with high probability, relying solely on the model’s intrinsic knowledge. This signifies a dependence on the parameter , such that . In contrast, reasoning with retrieval frameworks (Sun et al., 2024), posits the availability of a comprehensive knowledge base , and employs a retrieval model such that .
@@
While theoretically, this offers no benefit over RAG (Lewis et al., 2021) for generating answers with retrieved context, numerous studies (Ge et al., 2023; Bang et al., 2023) have shown that LLMs lack expertise in scientific domains and cannot construct a multi-step logical chain (Wei et al., 2023; Wang et al., 2023; Jiang et al., 2024) to connect a complex query with the retrieved incomplete information that cannot directly solve it. GIVE, in this context, provides additional guidance to LLMs, helping them to populate the expert knowledge towards the key components of the query by integrating both parametric and non-parametric knowledge.
@@ 3.2 Entity Group Construction
The set includes entities semantically akin to , and is added to to create the entity group :
@@ 3.3 Inner-group connections
Focusing on each in Section 3.2, we establish links between the queried entity and its semantically related concepts within its entity group. This aims to prompt the LLM to explore related concepts broadly, rather than focusing solely on the queried entity. The challenge of directly inducing relationships between two queried entities is addressed by identifying possible links between two sets of similar concepts. We employ the LLM to suggest relationships between the queried entity and each in-group concept. All such knowledge are appended to an affirmative knowledge set. Consider an entity group with 1 queried entity and additional KG concepts , for :
@@ 3.4.1 Potential Relations Induction
We first determine all possible relationships that could link any pair of nodes within the two groups. We evaluate two categories of potential relations: (1) Relations specified in the question. These are the vital connections required for the ultimate QA task. (2) KG relations present between these groups. Due to the semantic similarity of concepts within each group, existing cross-group KG connections may link other entity pairs across groups. Formally, for KG , entity groups and , we define the set of potential relations as:
@@ 3.4.2 Intermediate Node Group Discovery for Multi-step Reasoning
In scientific fields, directly establishing links between specific concepts is sometimes not feasible. For instance, to address a query regarding a drug’s effect on a disease, a logical approach is to associate these through (drug, compound, disease) links, forming the statement: "since a certain compound exists within the drug, and this compound alleviates particular diseases, it can be inferred that the drug may treat the disease." To support such multi-step reasoning process, GIVE investigates novel node groupings used as intermediate steps in reasoning. GIVE directs LLM to identify the most beneficial 2-hop paths between groups, helping with the question-answering task. Through the intermediate node of an optimal 2-hop path, GIVE forms an intermediate entity group using the process outlined in Section 3.2.
@@ 3.4.3 Structured expert knowledge guided reasoning
Finally, GIVE introduces associative thinking from these non-parametric evidence, including concept groups and all potential relations between them.
Relation Assignment Using External Evidence. If an edge exists in the external knowledge graph between two entities, we inherit the ground truth relationship from the original KG. When verbalizing such edges, we use a definite tense to convey the high certainty of this fact.
@@
Identifying Open Relations for Novel Connections To avoid the limitations of the knowledge base’s scope, GIVE also offers LLM the freedom to independently generate a short phrase describing the relation of each entity pair.
@@ 3.5 Progressive answer generation
From the reasoning process presented in Section 3.4.3, GIVE obtained three types of knowledge: (1) Affirmative knowledge set that contains all knowledge affirmed by LLMs’ parametric knowledge, including inner-group connections; the cross-group open connections, and the extrapolated connections. We refer to this knowledge set as . (2) Counterfactual knowledge set that contains all rejected potential cross-group connections, which is referred to as , (3) Ground-truth KG knowledge set . To prevent hallucination, we adopt a progressive manner to generate the final answer by first giving only the affirmative knowledge set. GIVE further directs LLM to refine this answer with the full context and the counter-factual knowledge set. The final answer is generated by providing all previous context and the ground-truth knowledge set. Given generator , and the retrieved knowledge sets described above, the answer generation process by GIVE is defined as:
where , , corresponds to the answer generation process that uses (1) only affirmative knowledge; (2) both affirmative knowledge and counterfactual knowledge; (3) the whole set of knowledge, which contains affirmative, counterfactual, and ground truth KG knowledge.
@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ B.1 Different ways of prompting
B.2 Number of seeded examples
B.3 Encoding model size
@@ C.1 Efficiency of GIVE
| Dataset | Avg Input Tokens | Avg Output Tokens | ||
| ToG | ToG | |||
| PubmedQA | 14518.5 | 12701.1 | 183.1 | 104.2 |
| BioASQ | 7970.3 | 7010.0 | 80.0 | 60.2 |
| ProcessBank | 19460.5 | 11995.6 | 232.5 | 100.5 |
| CSQA/10% ConceptNet | 5321.1 | 4934.8 | 34.9 | 23.1 |
| CSQA/50% ConceptNet | 7203.0 | 6704.1 | 45.2 | 34.8 |
| CSQA/100% ConceptNet | 7398.7 | 6679.7 | 46.3 | 36.6 |
@@
We further conduct experiments to compare the efficiency of GIVE with RAG (Lewis et al., 2021) and ToG (Sun et al., 2024), in terms of running time and context length, for every experiment setting we include in Section 4, the results are presented in Table 9. ALso, to conduct an in-depth analysis of token consumption, we compare the average number of input.output tokens for GIVE and ToG (Sun et al., 2024), as in Table 8.